OpenStudy (anonymous):

Terri is analyzing a circle, y2 + x2 = 36, and a linear function g(x). Will they intersect? y2 + x2 = 36 g(x) graph of the function y squared plus x squared equals 36 x g(x) -4 -4 -2 -2 2 2 Yes, at positive x-coordinates or zero. Yes, at negative x-coordinates or zero. Yes, at negative and positive x-coordinates or zero. No, they will not intersect.

OpenStudy (anonymous):

help please?

OpenStudy (anonymous):

I have the graph if you want it

OpenStudy (anonymous):



OpenStudy (anonymous):

i'M CONFUSED

OpenStudy (anonymous):

this line cuts the circle at two points denoted by cross.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

What would the answer be?

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

I think c

OpenStudy (anonymous):

correct but they do not intersect at zero.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

so

OpenStudy (anonymous):

Can you explain why its c please?

OpenStudy (anonymous):

if we denote g(x) by y then we see y=x  \[x^2+y^2=36,x^2+x^2=36,2x^2=36,x^2=\frac{ 36 }{ 2 }=18,x=3\sqrt{2},x=-3\sqrt{2}\] so it cuts both at negative and positive co-ordinate.